In a circle of radius 30 inches, a central angle of 35 will intersect the circle forming an arc of length _____.
1
Expert's answer
2020-05-18T17:19:33-0400
  1. an arc of length = L"\Delta"
  2. R = 30 inches
  3. "\alpha" = 35°.

"L\Delta=(πR/180)*\alpha",

"L\Delta=(π*R/180)*35=(π*30/180)*35=35π/6."